### v2

# Mind the Gap: The Divergence Between Human and LLM-Generated Tasks

Abstract
Humans constantly generate a diverse range of tasks guided by internal motivations. While generative agents powered by large language models (LLMs) aim to simulate this complex behavior, it remains uncertain whether they operate on similar cognitive principles. To address this, we conducted a task-generation experiment comparing human responses with those of an LLM agent (GPT-4o). We find that human task generation is consistently influenced by psychological drivers, including personal values (e.g., Openness to Change) and cognitive style. Even when these psychological drivers are explicitly provided to the LLM, it fails to reflect the corresponding behavioral patterns. They produce tasks that are markedly less social, less physical, and thematically biased toward abstraction. Interestingly, while the LLM’s tasks were perceived as more fun and novel, this highlights a disconnect between its linguistic proficiency and its capacity to generate human-like, embodied goals. We conclude that there is a core gap between the value-driven, embodied nature of human cognition and the statistical patterns of LLMs, highlighting the necessity of incorporating intrinsic motivation and physical grounding into the design of more human-aligned agents.
Figure 1: Overview of the study. (A) Conceptual framework. Human-generated tasks are typically driven by intrinsic motivation and grounded in embodied experience. In contrast, LLM-generated tasks are produced based on input prompts and their training data, which may result in a fundamental gap between the two. (B) Illustration of the thematic and embodiment gap of the tasks. Human-generated tasks tend to be more social and physically engaging, while LLM-generated tasks are less socially oriented and more abstract or cognitively focused.
Introduction
Humans can generate an infinite variety of goals to guide their behavior and enrich their daily lives, often infused with a distinct personal touch. This ability of autonomous task generation plays a central role in human cognition, shaping how individuals adapt to and interact with the world (Chu, Tenenbaum, and Schulz 2024; Molinaro and Collins 2023). Recent advances in large language models (LLMs) have enabled generative agents that simulate human activity in virtual worlds with striking autonomy (Park et al. 2023; Yang et al. 2024). These agents can initiate tasks, interact with others, and maintain activity over time, often producing goal-directed, human-like behavior. Yet the core question remains: does this simulation capture the deep cognitive mechanisms that drive human behavior, or is it merely a sophisticated mimicry of its surface patterns? Can LLMs, trained on vast corpora of text, truly replicate the diversity, personal flair, and intrinsic motivation that characterize human goal-setting?
Answering this question requires shifting from performance benchmarks to an analysis of the internal drivers behind task generation. Human goal-setting is not an unconstrained, probabilistic generation of action tokens. It is shaped by two core pillars of cognition.
The first is a system of value-driven motivation. Personal values are stable, trans-situational goals that serve as guiding principles in life (Sagiv et al. 2017; Schwartz 1992; Sagiv and Roccas 2021). They provide the core motive force for human actions, shaping what we find desirable or important. The value dimension of openness to change versus conservation, for example, reflects a deep-seated conflict between the pursuit of novelty and stimulation versus a preference for tradition and security (Schwartz 1992; Brosch et al. 2018), directly influencing creative and exploratory behaviors. These values shape not only what we do, but how we prioritize among competing possibilities.
Second, human goals are shaped by embodied experience, that is, by the sensorimotor constraints of the body and accumulated interactions with the physical and social world (Varela, Thompson, and Rosch 2017; Foglia and Wilson 2013; Xiang et al. 2023). Humans naturally understand objects not just by semantic labels, but by how they can be physically manipulated and used in context. This grounding shapes how we imagine, evaluate, and select possible actions in real-world settings.
These two factors are largely absent from current LLMs. While recent LLM-based generative agents can exhibit behaviorally rich outputs, it remains unclear whether they possess mechanisms for prioritizing goals based on internal needs or for evaluating the physical feasibility and affordances of objects. Although emerging approaches model goals as procedural programs (Davidson et al. 2024), the crucial influence of these deep psychological and embodied factors remains largely overlooked.
Research Objectives and Hypotheses
To systematically investigate these issues, we designed a novel text-based task generation paradigm that elicits unconstrained, intrinsically motivated responses from both humans and LLMs (Figure 1). Our approach proceeds in two steps.
First, we examine how human task generation reflects underlying personal factors, establishing a behavioral signature of autonomous task generation [^1] Our hypotheses were pre-registered at AsPredicted platform https://aspredicted.org/t3yb-wcd3.pdf. We hypothesize that if human task generation is indeed value-driven, then participants’ personal values (e.g., openness to change) should systematically predict key attributes of their generated tasks, such as their creativity and novelty. Furthermore, we investigate how individuals strategically adapt to environmental complexity by examining their cognitive style (Sagiv et al. 2010). We conceptualize cognitive style as the mechanism through which value-driven goals are pursued under varying cognitive loads. This aligns with dual-process theories distinguishing between deliberate, rule-based processing (systematic) and rapid, associative processing (intuitive) (Evans and Stanovich 2013; Sagiv et al. 2014). We reasoned that individuals relying on systematic processing would be more sensitive to the increased cognitive demand of complex environments, affecting the diversity and nature of their generated tasks.
Second, we compare this human baseline to the output of a widely used LLM (GPT-4o), conditioned on the same value. This allows us to evaluate whether such models can replicate the psychological signatures of autonomous goal-setting or whether fundamental differences remain, particularly in terms of motivational grounding and embodied realism.
Contributions
Our study yields the following key contributions:
  1. 1.
    Human Signatures of Autonomous Goal-Setting: We provide behavioral evidence that human task generation is systematically shaped by personal values and cognitive style, supporting a value-driven account of autonomous task generation.
  2. 2.
    LLM Outputs Lack Value-Driven and Embodied Signatures: Despite being prompted with individual human profiles (including personal values and cognitive styles), the LLM fails to exhibit core behavioral signatures of human goal generation.
    • No Induction of Human-Like Behavior: Conditioning on value-related inputs does not lead the LLM to generate behavior that mirrors human patterns of goal-setting. This suggests a lack of internal mechanisms for value prioritization or motivational grounding.
    • Thematic and Embodiment Gap: The LLM’s outputs show a strong thematic bias, favoring abstract over social or physical activities, and result in tasks perceived as more mentally demanding and less physically embodied than those generated by humans.
Related Work
Our research connects two domains: the psychology of autonomous goal-setting in humans and the challenge of simulating this behavior with LLMs, particularly concerning their psychological fidelity and embodied grounding.
Autonomous Goal-Setting and Individual Differences
Classic goal-setting theory has established how specific, difficult goals can enhance performance. However, much of human life is guided by autonomously generated goals in open-ended environments. Recent computational work has begun to model this process, for instance, by representing goals as programs in a domain-specific language (Davidson et al. 2024). While promising, these approaches often operate within constrained task spaces and do not account for the diverse psychological factors that drive human choice.
Human goals are not merely procedural; they are expressions of stable and diverse motivations. A large body of psychological research shows that personal values, trans-situational life principles like security, achievement, or benevolence, systematically guide attitudes and behavior (Sagiv and Roccas 2021; Sagiv et al. 2017; Kasof et al. 2007). Similarly, cognitive styles, such as the preference for systematic versus intuitive thinking, shape how individuals approach problems and adapt to environmental complexity (Evans and Stanovich 2013; Sagiv et al. 2014). We incorporate both factors to define a human behavioral baseline for autonomous goal generation.
LLMs as Cognitive Simulators: Capabilities and Limitations
Recent advances have positioned LLMs as powerful tools for simulating human cognition and behavior. Studies have demonstrated that LLM-based “generative agents” can navigate complex environments, autonomously generate plans, and engage in social interactions that exhibit emergent dynamics (Park et al. 2023; Yang et al. 2025). These capabilities have led to growing interest in using LLMs as cognitive models to explore how humans think, decide, and act.
Despite these capabilities, researchers have identified significant gaps in the psychological plausibility of these models. While LLMs can mimic surface-level reasoning and planning, they often lack deeper cognitive mechanisms such as emotion, causality, physical dynamics and social cognition (Hu, Sosa, and Ullman 2025; Binz and Schulz 2023; Ullman 2024). Moreover, because they are trained purely on text, LLMs lack the sensorimotor grounding that supports embodied goal formation in humans. Embodied cognition theories emphasize that human goals and reasoning are deeply shaped by our physical experiences and social embeddedness (Varela, Thompson, and Rosch 2017). This suggests that generating goals rooted in physical interaction and embodied experience is a critical test case for the fidelity of LLM simulators—a test we conduct in this study.
Methods
We conducted two experiments to identify the unique signatures of human task generation by contrasting it with the output of a large language model (LLM). Experiment 1 elicited tasks from both human participants and an LLM agent. For the human baseline, we measured core psychological constructs hypothesized to drive autonomous behavior: stable motivational drivers (personal values, via PVQ21, Schwartz 2021), cognitive strategies for navigating complexity (thinking style, via TWS, Sagiv et al. 2010), along with emotional states as controls. Experiment 2 then involved an independent sample of human raters who evaluated key attributes of the generated tasks, enabling a direct comparison between human and AI outputs.
Figure 2: Experimental interface and procedure. (A) Text-based task generation interface. Participants were asked to generate tasks using a given set of room items. They were instructed to report the task name, required items, detailed setup, goals, and scoring rules. (B) Experimental conditions. Participants were randomly assigned to one of four room scenarios, varying in environmental complexity (high vs. low) and social context (presence vs. absence of other people), sampled from a virtual simulation platform. (C) Task evaluation phase. Independent raters assessed both human- and LLM-generated tasks across multiple dimensions, including Fun, Novelty, Mental Demand, and Physical Demand.
Experiment 1: Human and AI Task Generation
Participants
A total of 180 participants were recruited via the Prolific platform (https://www.prolific.com/). 4 participants were excluded for meaningless task response, resulted in 176 participants in total (110 female, 66 male; age range: 18-65). All participants were fluent in English and provided informed consent prior to the experiment. Participants were paid approximately £6.4/h for their participation.
Stimuli and Design
The experiment employed a 2 (environmental context: high vs. low complexity) $\times$ 2 (social context: social vs. non-social) between-subjects factorial design (Figure 2B)  [^2] Experiment link: https://value-task-collect.pages.dev/consent. The environmental context was manipulated to vary the creative constraints. In the high-complexity condition, participants were presented with a list of 32 common living room items. To increase the cognitive demand for task generation in the low-complexity condition, we removed seven items with high entertainment affordances (e.g., Dice, Basketball), leaving 25 common items (e.g., Pens, Chairs).
The social context was manipulated to examine the influence of social motivation and embodied understanding. The social condition included a “person” icon in the item list to enable interaction-based goals. The non-social condition replaced “person” with “mannequin” to present a human-like form devoid of social agency. This manipulation isolates goal generation cued by social presence from that cued merely by a humanoid shape.
All object stimuli were selected from the room asset library of the virtual simulation based on Unreal Engine 5 platform  [^3] https://www.unrealengine.com/en-US/unreal-engine-5 for ecological validity.
Psychological Measures
Personal value priorities were measured with the PVQ21 (Schwartz 2021). Following established procedures, we computed scores for two orthogonal dimensions: openness to change (openness vs. conservation) and self-interest (self-enhancement vs. self-transcendence). Thinking style was measured with the Thinking and Working Style (TWS) Questionnaire (Sagiv et al. 2010), which places individuals on a continuum from intuitive to systematic thinking. We administered the sociability facet of the Big Five Inventory–2 (BFI-2) (Soto and John 2017) as a targeted measure of the propensity for social interaction. The emotional valence and arousal for the day were also recorded to control participants’ short-term mental states.
Procedure
After completing the psychological scales, participants were randomly assigned to one of the four conditions. They were shown a list of objects and instructed: “Please imagine that you are in a real room with several items in it. You need to use these items to pass the time.” For each task, participants provided a title, a list of used objects, a description, task setup, goal, and scoring system (Figure 2A). They reported the mental effort and perceived difficulty for each task generated. Participants were required to generate at least three tasks before choosing to conclude the experiment.
LLM Agent Implementation
We implemented two GPT-4o agent conditions to test whether the model’s deviation from human goal-setting stems from a simple information deficit or a fundamental mechanism differences. In the raw condition, the model received only the task prompt, serving as a baseline for its default behavior. In the matched condition, each agent was given the full psychological profile of a corresponding human participant, including values and cognitive style. This condition directly tested whether the model could translate these psychological constructs into value-driven behavior when provided as semantic data, or whether its lack of internal motivation and embodied grounding imposes a deeper constraint. Both agents operated at a temperature of 1.0 to encourage diverse outputs.
Experiment 2: Task Attribute Evaluation
Participants
77 participants were recruited via an online platform. Two were excluded for random responding, leaving 75 evaluators (35 male, 30 female; age range: 18-31). They were compensated for their time.
Stimuli and Procedure
The stimulus pool for evaluation comprised two sets of tasks. To capture the full spectrum of human creativity, we included all 575 valid tasks generated by human participants. To create a focused and stringent comparison, we compared this to the output from the matched GPT condition, including the first task generated for each of the 176 unique participant profiles. This ensures a representative, non-redundant sample of the model’s performance when conditioned on individual data. Tasks from the raw GPT agent were reserved for computational analyses (e.g., topic modeling) but excluded from human rating to streamline the evaluation process. Each of the 75 evaluators then rated a randomized subset of 50 tasks, yielding approximately five independent ratings per task.
Evaluation Metrics
Evaluators rated each task on a 0-10 scale across several dimensions selected to capture creativity (Fun, Novelty), and embodiment (Mental Demand, Physical Demand, Body Part Engagement). These metrics provide a quantitative basis for comparing the qualitative nature of human- and AI-generated goals (Figure 2C).
We assessed inter-rater reliability using the Intraclass Correlation Coefficient (ICC, Koo and Li 2016) based on a two-way random-effects model. Given the variable number of raters per item, we fitted a linear mixed-effects model to estimate variance components and calculated ICC(2,k) using the average number of raters (k). The ICCs for task dimensions rated by participants ranged from 0.58 to 0.90, indicating moderate to good reliability (Koo and Li 2016). Given that dimensions such as Fun involve considerable subjectivity, these values suggest that the measurements were acceptably reliable.
Data Analysis
Task Content Analysis
To identify emergent themes in the generated tasks, we performed topic modeling using BERTopic (Grootendorst 2022) on the full corpus of 1718 tasks (575 human tasks, 575 matched GPT tasks, and 568 raw GPT tasks). To measure the creativity of each participant’s output, we also computed their task diversity score, defined as the mean cosine distance between the sentence-transformer embeddings (Reimers and Gurevych 2019) of the tasks they generated.
Statistical Modeling
Figure 3: Personal values and cognitive style shape human goal generation. (A) Regression coefficients of the key predictors on task attributes (Novelty, Fun, and Task Diversity). (B) Openness to Change values predict task fun and novelty. (C) Cognitive style (TWS) interacts with environment to predict task diversity. Bars indicate 95% confidence intervals. * indicates $p<0.05$, ** indicates $p<0.01$.
Human Baseline Analysis.
To test our hypotheses about the link between psychological traits and task attributes, we specified a series of linear mixed-effects models (LMMs, Brauer and Curtin 2018) to predict key dependent variables: the rated attributes of Fun, Novelty, and their task content diversity. We first specified a baseline model including only demographic variables and random intercepts for participants. Then we specified a main-effects model by adding experimental conditions and psychological predictors like personal values and thinking style. Finally, we additionally included the pre-specified TWS × Environmental Context interaction into the model. The models were compared using a likelihood ratio test (LRT, Moreira 2003) to determine if including the psychological predictors significantly improved model fit.
Human-LLM Comparison.
We used chi-square tests to compare the distribution of task topics and the frequency of multiplayer task generation. We used Mann-Whitney U tests (McKnight and Najab 2010) with false discovery rate (FDR) correction to compare the rated attributes of human- and AI-generated tasks.
Results
Our analyses proceed in two stages. First, we establish a behavioral signature of human goal generation, demonstrating its systematic link to personal values and environmental context. Second, we contrast this human baseline against the output of LLM agents to reveal fundamental mechanism gaps.
The Signature of Human Goal Generation: Value-Driven and Environmentally Sensitive
To establish a human baseline, we first tested how psychological traits predicted the goals people generated. Linear mixed-effects models confirmed that task generation was systematically linked to stable personal values. The value of Openness to Change significantly predicted higher ratings of both task novelty ($b=0.152$, $t=2.406$, $p=0.017$) and fun ($b=0.146$, $t=2.316$, $p=0.022$). In contrast, the self-interest value dimension showed no such relationship. This finding supports the core hypothesis that stable motivational principles guide autonomous goal-setting.
Behavior was not only driven by stable values but was also adapted to the environment. We found that the creativity of generated tasks depended on a significant interaction between an individual’s cognitive style (TWS) and the complexity of the environment. This pattern appeared in both subjective ratings of task fun (interaction of TWS and Environmental Context: b=0.140b=-0.140, t=2.640t=-2.640, $p=0.009$), and in objective measures of task diversity (b=0.215b=-0.215, t=2.809t=-2.809, $p=0.006$). Specifically, individuals with a more systematic thinking style generated more fun tasks in the simple environment, whereas those with an intuitive style thrived in the complex one. This suggests people flexibly deploy different cognitive strategies to match environmental demands.
Figure 4: Human and GPT generated systematically different goals. (A) Thematic distribution of human and LLM tasks. 1718 tasks generated by human and GPT were clustered into 12 topics and 3 themes: “Physical & Sports Activities”, “Relaxation & Household Activities”, and “Mental & Artistic Activities”. (B) LLMs exhibit a lower propensity for social task generation. (C) LLM tasks are perceived as more mentally and less physically demanding. Black dots stands for the overall difficulty of tasks. Orange and green dots stands for mental and physical load respectively. (D) Engagements of different body parts.
The Human-LLM Gap: A Disembodied and Asocial Simulation
Having established a human behavioral signature, we next compared it to the output of LLM agents to identify structural differences. The comparison revealed systematic differences in task content and structure.
Thematic Differences in Task Generation
We directly compared the distribution of task topics between human participants and LLM agents. Topic modeling identified three dominant themes across the dataset: Physical & Sports, Relaxation & Household, and Mental & Artistic. The distribution of tasks across these themes differed significantly between humans and the LLM agents ($\chi^{2}(4)=581.45$, $p<0.001$). Human participants generated a balanced mix of activities. In contrast, LLMs showed a bias toward abstract tasks like music and writing (Figure 4A).
This bias was most evident in the Matched GPT condition, where “Mental & Artistic” tasks accounted for 74% of its output. For example, music-related tasks (“kalimba music”) were rare in human responses (1%) but common for both Raw GPT (13%) and Matched GPT (11%). This focus on the abstract resulted in the near-total neglect of the “Relaxation & Household Activities” theme by the models (3% for Raw GPT, 5% for Matched GPT), which was a major component of human-generated goals (32%). Consequently, common embodied tasks involving using everyday objects like hangers and treadmills were also almost absent from the LLM’s responses.
Limited Social Interaction in LLM Tasks
Beyond thematic content, the models generated significantly fewer social tasks than humans in settings where another person was present ($\chi^{2}(2)=93.19$, $p<0.001$). In these scenarios, 58% of human-generated tasks were multiplayer, with 83% of participants creating at least one social goal. Humans proposed a range of interactive tasks, from simple conversation games like “Guess Who I Am” to competitive games like “Pillow Fight”. In contrast, multiplayer tasks from the Raw GPT and Matched GPT were far less common, accounting for only 29% and 24% of their outputs, respectively (Figure 4B).
A Disembodiment Gap in Perceived Demands
The thematic differences between human and LLM tasks were reflected in their perceived mental and physical demands (Figure 4C). Tasks generated by the Matched GPT were rated as significantly more mentally demanding (z=11.64z=-11.64, FDR-corrected $p<0.001$) and less physically demanding (z=8.40z=-8.40, FDR-corrected $p<0.001$) than those generated by humans. This aligns with the LLM’s focus on abstract, “Mental & Artistic” activities.
This mental-physical divide was also evident in the specific body parts required for task execution (Figure 4D). GPT tasks demanded greater brain engagement (z=11.40z=-11.40, FDR-corrected $p<0.001$) but significantly less involvement of the arms (z=5.43z=-5.43, FDR-corrected $p<0.001$) and lower body (z=7.93z=-7.93, FDR-corrected $p<0.001$). Engagement of the eyes, ears, and hands did not differ between the two groups.
The Paradox of Ungrounded Creativity
Despite their disconnection from physical and social experience, Matched GPT tasks were rated as more novel (z=5.63z=-5.63, FDR-corrected $p<0.001$) and more fun (z=2.73z=-2.73, FDR-corrected $p=0.010$) than those created by humans (Figure 5). This highlights a key tradeoff: LLMs can generate text that seems creative and engaging, but they do so without grounding in embodied or socially situated experience.
Figure 5: LLMs generate tasks rated as more fun and novel.
Discussion
This study examined the cognitive basis of autonomous goal generation by comparing human behavior with large language models (LLMs). Our findings show that human-generated goals are systematically shaped by personal values and are adapted to environmental complexity. In contrast, LLMs, even when prompted with personal profiles, exhibit fundamental and predictable biases, revealing a significant gap between their text-based simulation and the grounded nature of human cognition.
Drivers of Human Goal Generation
Individual differences shaped goal content. People high in openness to change generated more novel and fun tasks, consistent with prior work linking openness to exploratory behavior and intrinsic motivation (Deci and Ryan 2013; McCrae and Costa Jr 1997; Sagiv and Roccas 2021). Cognitive style also mattered: systematic thinkers generated less diverse tasks in complex environments, possibly due to higher cognitive load. These results align with previous creativity and decision-making theories, which suggest that systematic thinkers prefer structured problem-solving, while intuitive thinkers are more adaptable in diverse settings (Evans and Stanovich 2013; Sagiv et al. 2014).
Explaining the Human-LLM Gap
The systematic differences we observed point to two core mismatches.
First is the embodiment gap. Humans generated tasks involving physical interaction with common objects (e.g., hangers, pillows), reflecting sensorimotor knowledge of object use. LLMs rarely produced such tasks. Their outputs were rated as less physically demanding and less involving of limbs. This supports the view that LLMs rely on text-based associations and lack access to intuitive physics or affordance perception (Lake et al. 2017).
Second is the value and motivation gap. Humans often proposed social and prosocial tasks, consistent with basic motivational drives such as affiliation and well-being. In contrast, LLMs showed a lower propensity for social engagement. Their strong bias toward abstract, mental tasks (e.g., music, writing) and away from relaxation or household chores points to a lack of an internal, value-driven system. Crucially, providing the model with user’s psychological profile did not fix this. This finding suggests that simply describing values in a prompt is not equivalent to having an internal value systems that regulates task selection.
The Paradox of Aligned Creativity
LLM-generated tasks were rated as more novel and fun. This highlights the unique nature of LLM creativity. Unconstrained by physical feasibility, LLMs excel at combinatorial creativity, drawing on their vast training data to produce imaginative textual descriptions.
This natural tendency is likely amplified by current alignment methods like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022; Bai et al. 2022). Such methods optimize for outputs that human raters find immediately helpful, engaging, or interesting. This process inherently discourages the generation of mundane, routine, or repetitive actions like resting or cleaning. Although these activities are central to daily life, they are less likely to receive high ratings and are therefore filtered out. The result is a model fine-tuned to produce “ungrounded” but highly-rated novelties, which explains both its high creativity scores and its thematic detachment from the reality of human experience.
Future Directions and Limitations
Our findings have two main implications. For AI research, they suggest that creating truly human-like autonomous agents requires more than scaling up language models. Addressing the embodiment and motivation gaps may require integrating world models, intrinsic reward systems, or sensorimotor learning (Matsuo et al. 2022). For cognitive science, our task-based paradigm offers a scalable, flexible tool to study goal-setting in naturalistic contexts, beyond traditional lab constraints. It enables new comparisons between human cognition and machine outputs.
Several limitations of our study point toward promising future directions. First, the text-based task generation paradigm may not fully capture the richness of human goal-setting, which often involves multimodal, interactive, and context-sensitive elements. Future work could incorporate interactive 3D environments to enhance ecological validity. Second, due to the high cost of human evaluation, we restricted our comparisons to a single LLM architecture (GPT-4o). Expanding the scope to include a broader range of models with different sizes and architectures could offer a more comprehensive understanding of how different model properties influence goal generation. Finally, the diversity and complexity of human-generated goals highlight the limitations of current computational frameworks, and call for new approaches capable of modeling open-ended, value-guided behavior.
Acknowledgments
This work was partially supported by the National Science Foundation of China under grant No. 32441106.